A standing wave on a string of length L = 3 m is described by the following equation: y(x,t) = 0.08 sin⁡(πx) cos(300πt). The fundamental frequency, f1, is:


1
Expert's answer
2021-05-20T10:09:41-0400
"f_1=\frac{\omega}{2\pi}\\f_1=\frac{300\pi}{2\pi}=150\ Hz"